UT 20131213 LUCI2 Nighttime Commissioning

Observer: JHeidt, DThompson
Telescope Operator: GBechetti
LBTO Support:

Plan for the night

Weather is bad (fog + some snow), unlikely that we open => closed-dome tests

- Tests G150 stabilization
- Reproducability tests gratings
- Focus determination N1.8 camera
- Linearization map

Summary

- High humidity all night
- Repro-tests finished
- Persistence/crosstalk data taken

- G150 not very stable => will not be considered for the rest of COM
- Scripts stopped several times without error => AI Alexander

Preparations

- Alexander will come tomorrow, got stuck in London due to a delayed flight
- Jochens arrival was late due to nail in tire and problems with snow chains

Details

Files are luci2.20131212.xxxx

Login
Start GEIRS gui, take image...OK
Start LMC
Restart GEIRS Readout, Readout Manager, Scheduler
Start Obserer gui
Open LUCI2 and READ2 guis
Take image...OK

#0001# 0002 test exposures (forgot to check manual save)

Telescope @ horizon, lights on in enclosure

21:00 UT on the 12th
Dark_RN.xml script already loaded
reset, commit 6 items

#0003 #0022 1x2.7399s LIR
#0023 #0042 1x6.8515s MER
#0043 #0046 1x60s MER
#0047 #0050 1x300s MER
#0051 #0054 1x600s MER

Walter did tests on G150 stabilization, FM1 & FM4, checked the alignment of FoV,
and updated the focus positions for all filters

0:20 Sunset: 100% humidity, fog and occasionally snow showers, T=-1deg, dome stays closed

Grating reproducability tests

03:00 ROTANGLE=210
03:04 Grating tilt repro: G150, didn't stabilize at 2.22mue (4th exp), #56-58
spectra look crazy shifted by a large amount
03:09 Redo, went through, #59-64
03:16 Throughslit image (0.5" slit), looks OK, #65
03:20 AGw in beam, Geno took it out, Grating tilt G200, spectra ok, #66-71
03:26 grat tilt G210, spectra ok, #72-77
03:33 grat select: #78-86
03:50 Did a few checks with G150, lambda-voltage seems to be off by 1V!!!,
should be 21.7mue but is actually the cutoff of the KS-filter at 2.3mue, #87-88

03:55 ROTANGLE=300
03:56 grat tilt g150, stopped at 2.22mue, #89-91
04:01 Redo, this time ok, #92-97
04:05 grat tilt G200, #98-103
04:08 grat tilt G210, #104-109
04:11 grat select, #110-118

04:17 ROTANGLE=120
04:18 grat tilt G150, didn't want to stabilize at all
04:23 grat tilt g200, stabilizes, but no data taken, stop/start ROM, scheduler, reini IM
Fw2 got stuck, recovered (Note: when coming from Pos15-> Pos1, reference switch doesn't
automatically go active, when coming from Pos2->Pos1 it does!
04:48 retry grat tilt G200, OK, #119-124 04:52 retry grat tilt G150, FW2 got stuck again, recover, restart script, OK #125-130
05:07 grat tilt G210, FW2 got stuck again, recovered, restart script, #131-136
05:20 grat select, FW2 got stuck once more (moving from K->HKspec)!!! #137-139
recover, set motion direction to negative, restart script, FW2 got stuck again (HKspec->Ks),
recover, restart script,#140-143, recover, set motion direction to positive, restart script,
FW2 got stuck again, recovered but stopped

06:11 ROTANGLE=300
06:12 Go back to ROTANGLE=300 to see if FW2 problem persists, start grat select, OK,#144-152
Tilt of G150 is OK now, don't understand why, no problem with FW2 nor G150

06:21 ROTANGLE=30
06:22 grat select, another test for FW2/G150, OK, #153-161

06:28 ROTANGLE=120
06:29 grat select again, see what happens, #162-170
Who understand this, everything OK!!!
06:35 grat tilt G150, didn't stabilize after 2nd exp., #171-172
06:39 retry grat tilt G150, OK, #173-178
06:44 grat tilt G200, OK, #179-184
06:47 grat tilt G210, OK, #185-190

06:49 ROTANGLE=30
06:50 grat tilt G210, #191-196
06:52 grat tilt G200, #197-202
06:55 grat tilt G150, #203-208
06:59 grat select, #209-218

OK, G150 stabilization test and grating repro-tests done

Focus determination N1.8-camera

New value for K set by Walter to -285, needs to be rechecked for zJHK.
Expected values are z=-259, J=-264, H=-278, K=-285
07:17 K, start = -310, step=5, nsteps=11, no lamp, 3sec, #218-228
07:21 H, start = -305, step=5, nsteps=11, lamp3, 3sec, #229-239
07:24 J, start = -290, step=5, nsteps=11, lamp3, 3sec, #240-250
07:27 z, start = -285, step=5, nsteps=11, lamp2, 3sec, #251-261
Seems Ok for K, but not for zJH, redo zJH with adjacent values to have full set of focus data
07:31 z, start = -335, step=5, nsteps=11, lamp2, 3sec, #262-272
07:35 J, start = -340, step=5, nsteps=11, lamp3, 3sec, #273-283
07:38 H, start = -355, step=5, nsteps=11, lamp3, 3sec, #284-294

08:00 Partly clear now, but humidity still 100%, Temp has dropped to -6deg.

Linearity tests

Running OB: Linearize.xml
Takes data to determine linearization coefficients
Pairs of images at each DIT, running from 2.75 to 100.0 seconds and back down,
with a 2.75s pair between each step. J-high filter and Halo3 lamp.
#0295-0362

Execution stopped at the 2.75s exposures after the descending 90s
Had to stop/start lots of stuff, including killing java processes and recover FW2.

Starting the whole script again.
UT 09:31 Error committing instrument...need to initialize all
UT 09:35 start Linearize.xml OB
active item not updating...click refresh works

#0363-0394 script again stopped with no error messages.

Full shutdown and restart of SW in LMC
Initialize all units

UT09:59 start Linearize.xml OB
active item not updating

#0395-0424 script again stopped
regular messages: "Connection refused to host 192.168.60.12

Persistence tests

#0425 test image through diagonal mask J filter, Halo3
#0426 1x2.7399s Nexp=10 INT/LIR
#0427 1x2.75s Nexp=10 INT/LIR
#0428 1x10.0s Nexp=10 INT/LIR
#0429 1x100.0s Nexp=1 NORM/LIR
#0430 1x300.0s Nexp=1 NORM/LIR

blind filters, lamp off (but diagonal mask)

#0431 1x300.0s to show persistence
#0432 1x2.7399s Nexp=10 INT/LIR
#0433 1x2.75s Nexp=10 INT/LIR
#0434 1x10.0s Nexp=10 INT/LIR #0435 1x100.0s Nexp=1 INT/LIR (ye
s, forgot to switch to normal)
#0436 1x300.0s Nexp=1 NORM/LIR

#0437-0441 1x300.0s to trace out persistence some more.

-- JochenHeidt - 13 Dec 2013
Topic revision: r8 - 13 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback